OpenStudy (rebeccapink6398):

The area of a rectangle is 61.64 square feet and the length is 9.2 feet. What is the width of this rectangle?

jimthompson5910 (jim_thompson5910):

A = L*W 61.64 = 9.2*W solve for W to get W = ???

OpenStudy (accidentalaichan):

61. 64 / 9.2 what do you get?

OpenStudy (accidentalaichan):

W = A/ L

OpenStudy (accidentalaichan):

sigh W = 6.7